You need to order carpet for a room that has a length that is twice its width. If the area of the room is 512, find the width of the room.
Answer by cleomenius(959)

W = x
L = 2x
2x^2 = 512
x^2 = 256
x = 16 this is the width
2*16 = 32 this is the length
32 * 16 = 512.
This does check.
Cleomenius.